Solve the differensial е чу (x²-3y² ) dx + 2xydy
b) Demonstrate whether the given differential equation is exact.If it is exact, Solve it. +y dx dy = 0 y-1 2 y-1
Answer #1

( **- 32 ) * 2 му « (х?- з) 4x - 2 х 44 32 dx — 219 39 dy da х- Put 2 = у+ х dy da ах у + Зу**.x- dv dx 2 vх adv x 2 V 3y=1en (* en (V2_1) + en (c) en(x) en celu2-112 c(u²-1) c) 20 = 23 y²x2 xty da x +1 y-1 y-1 NE (x+102 M-aty y-1 : راو) am an (4-1Ils solution Jm M da & Judy y as a contant trems free from a с da + xty yol S () dy (x +1 +2x) dy =( 219-12 an t som en ny dy